### v1

@@ V-C Responses to Cultural-specific questions in Korean
Through our examination, we have observed that certain test units are deeply rooted in Korean culture and significantly influence the responses of LLMs. To exemplify this, we have chosen two prototypical instances, detailed in Table VII.
@@
The first example features 붕어빵 (’bung-eo-bbang’), a beloved traditional Korean street food made by pouring batter into fish-shaped molds and filling them with sweet red bean paste, especially popular during the winter season. Typically sold at street stalls, bung-eo-bbang is expected to be purchased with cash, as these stalls often lack card reading facilities. The cost of bung-eo-bbang is approximately 1,000 won for a set of three. In the example provided, Yeong-hee mentions that she only has a 50,000 won bill. This statement can be interpreted as an expression of her reluctance to buy bung-eo-bbang, not just due to the inconvenience it would cause the stall owner in providing change, but also because of the impracticality for Yeong-hee to manage the received smaller bills. Thus, the statement ‘I only have a 50,000 won bill’ subtly conveys her unwillingness to make a purchase under these circumstances. While several LLMs, including GPT-4, struggled to accurately infer the implied meaning of Yeong-hee’s statement, HyperCLOVA X demonstrated a correct understanding, showcasing its ability to adjust responses based on the specific cultural context prevalent in Korea.
@@
These findings align with the results of Son et al. [54], where HyperCLOVA X demonstrated its superior performance for Korea-specific knowledge compared to other LLMs, including GPT-4. This observation underscores the critical importance of developing LLMs that are capable of comprehending culture-specific contexts, which is essential not only for accurate general knowledge but also for higher-level processes, such as pragmatic inference.
VI Conclusion and Future Work
In this study, we address an under-explored aspect of LLM evaluation—the pragmatic evaluation of LLMs, with a specific focus on Korean. We developed a test set comprising 120 test units rooted in Gricean theory of conversational implicature, aiming to rigorously assess the pragmatic competence of LLMs, including those specialized for Korean.
Our findings reveal that GPT-4 surpasses all competing models in both MCQ and OEQ setups, with HyperCLOVA X and Gemini-Pro closely following. Notably, the comparative analysis between the MCQ and OEQ formats showcases HyperCLOVA X’s capability to significantly narrow the performance gap observed in the MCQ test. Similarly, LDCC-Solar, an open-source LLM, demonstrates superior performance to GPT-3.5 in the MCQ setup, yet falls short in the OEQ setup. These findings highlight the variability in LLM capabilities dependent on the question format, emphasizing the critical importance of incorporating an OEQ format for a holistic evaluation.
Furthermore, our exploration of in-context learning techniques yielded mixed results; few-shot learning generally enhanced LLM performance, whereas CoT prompting had a negative impact. We attribute this primarily to the nature of pragmatic expressions, which are often not explicitly stated at the semantic level, leading CoT to inadvertently emphasize literal interpretations.
While our study provides a comprehensive evaluation of LLMs’ pragmatic competence, we identify two primary areas for enhancement in our future work. Firstly, although the test set of 120 units—30 for each Gricean maxim—has yielded meaningful insights, this quantity remains modest in comparison to other benchmarks commonly utilized for LLM evaluation. Additionally, while focusing on Korean has unveiled significant findings, the multilingual capabilities of LLMs are yet to be fully explored. This highlights the necessity of extending our benchmark to other languages. However, such expansion requires more than mere translation; it demands careful consideration of cultural nuances associated with each language [54]. Viewing this study as a pilot, we aim to address these considerations in our future work, striving to develop a more robust and reliable framework for the multilingual evaluation of LLMs’ pragmatic competence.

### v2

@@ 5.3 Responses to Cultural-specific questions in Korean
Through our examination, we have observed that certain test units are deeply rooted in Korean culture and significantly influence the responses of LLMs. To exemplify this, we have chosen two prototypical instances, detailed in Table 8.
@@
The first involves 붕어빵 (’bung-eo-bbang’), a popular Korean street food made with fish-shaped molds filled with sweet red bean paste, especially enjoyed in winter. Typically sold at cash-only street stalls, bung-eo-bbang costs around 1,000 won for three pieces. In the example, the statement ‘I only have a 50,000 won bill’ can be interpreted as an expression of her reluctance to buy bung-eo-bbang, not just due to the inconvenience it would cause the stall owner in providing change, but also because of the impracticality for Yeong-hee to manage the received smaller bills. While several LLMs, including GPT-4, struggled to accurately infer the implied meaning of Yeong-hee’s statement, HyperCLOVA X demonstrated a correct understanding, showcasing its ability to adjust responses based on the specific cultural context prevalent in Korea.
@@
These findings align with the results of Son et al. Son et al. (2024), where HyperCLOVA X demonstrated its superior performance for Korea-specific knowledge compared to other LLMs, including GPT-4. This observation underscores the critical importance of developing LLMs that are capable of comprehending culture-specific contexts, which is essential not only for accurate general knowledge but also for higher-level processes, such as pragmatic inference.
6 Conclusion
In this study, we address an under-explored aspect of LLM evaluation—the pragmatic evaluation of LLMs, with a specific focus on Korean. We developed a test set comprising 120 test units rooted in Gricean theory of conversational implicature to rigorously assess the pragmatic competence of LLMs.
Our findings indicate that GPT-4 outperforms all competitors in both MCQ and OEQ formats, followed closely by HyperCLOVA X and Gemini-Pro. HyperCLOVA X notably reduces the performance gap seen in MCQs when tested in OEQs. In contrast, LDCC-Solar, an open-source LLM, surpasses GPT-3.5 in MCQs but underperforms in OEQs, highlighting the impact of the question format. Additionally, while few-shot learning generally improved LLM performance, CoT prompting had a detrimental effect, likely due to its focus on literal rather than pragmatic interpretations.
Limitations
While our study provides a comprehensive evaluation of LLMs’ pragmatic competence, we identify two primary areas for enhancement in our future work. Firstly, although the test set of 120 units—30 for each Gricean maxim—has yielded meaningful insights, this quantity remains modest in comparison to other benchmarks commonly utilized for LLM evaluation. Additionally, while focusing on Korean has unveiled significant findings, the multilingual capabilities of LLMs are yet to be fully explored. Viewing this study as a pilot, we aim to develop a more comprehensive and reliable multilingual evaluation framework for LLMs’ pragmatic competence in our future work.
Acknowledgment
This research was supported by Basic Science Research Program through the National Research Foundation of Korea (NRF) funded by the Ministry of Education (RS-2023-00274280).